Musich Corporation has an activity-based costing system with three activity cost pools--Machining, Setting Up, and Other. The companys overhead costs, which consist of equipment depreciation and indirect labor, have been allocated to the cost pools already and are provided in the table below Machining Setting Up Other Tota $8800 S 47,200 S 23,200 S 79,200 Equipment depreciation Indirect labor Total 3600 2600 3800 10,000 $ 12,400 S 49,800 S 27.000 $ 89,200 Costs in the Machining cost pool are assigned to products based on machine-hours (MHs) and costs in the Setting Up cost pool are assigned to products based on the number of batches. Costs in the Other cost pool are not assigned to products. Data concerning the two products and the companys costs appear below: MHs Batches Product Z3 Product T1 Total 5700 5900 11,600 600 1200 Product Z3 Product T1 Sales (total) Direct materials (total) Direct labor (total) $ 224,800$ 252,500 $82,500 $ 97,000 $ 109,400$103,700FIND:

machining activity rate

setting up activity rate

amount of Overhead applied to Z3

amount of Overhead applied to T1

Product margin Z3

Product margin T1
Answer #1

1) Machining activity rate = 12400/11600 = 1.07 per machine hour

2) Setting up activity rate = 49800/1800 = $27.67 per setup

3) Overhead applied to Z3 = (12400/11600*5700+49800*600/1800) = $22693

4) Overhead applied to T1 = (12400/11600*5900+49800*1200/1800) = $39507

5) Product margin Z3 = 224800-82500-109400-22693 = $10207

6) Product margin T1 = 252500-97000-103700-39507 = $12293